OpenStudy (anonymous):

find the simple interest...principle $2000 Rate 10 and 1/2%....3 and 1/2 years

OpenStudy (anonymous):

Simple Interest, I = PRT P=principle, R=rate(%), T=time(yrs) I=2000(10.5/100)(3.5) I=735